OpenStudy (anonymous):

Analyze the function f(x) = sec 2x. Include: - Domain and range - Period and Amplitude - Two Vertical Asymptotes

OpenStudy (eyust707):

Domain- What x's are you allowed to put in Range - What y-vaules can the function spit out

OpenStudy (eyust707):

We are allowed to put in any x value

OpenStudy (anonymous):

so for domain & function what would i put?

OpenStudy (eyust707):

Domian- Any value of x... All real numbers

OpenStudy (ghazi):

there will be no amplitude since there is no maximum value of y so amplitude= none, period = 2*pi/2=pi

OpenStudy (eyust707):

oh wait the range is acutally all realls as well

OpenStudy (ghazi):

domain = any value of x

OpenStudy (ghazi):

vertical asymptote are x= pi/2, -pi/2

OpenStudy (anonymous):

omg thanks so much :) but range will be all real #s?

OpenStudy (anonymous):

@eyust707

OpenStudy (eyust707):

yea

OpenStudy (eyust707):

here is the graph http://www.wolframalpha.com/input/?i=f%28x%29+%3D+sec+2x

OpenStudy (anonymous):

thanks sweetie :)

OpenStudy (eyust707):

as you can see there are y values from -infinity to infinity

OpenStudy (ghazi):

yea it's correct range is all real numbers

OpenStudy (anonymous):

I still don't understand exactly how to find the period can you please help break it down.